OpenStudy (anonymous):

find the solution to the system of equations below, start by adding the equations. -2x + 2y = -8 2x + y = 5

pooja195 (pooja195):

-2+2=?

OpenStudy (anonymous):

0?

pooja195 (pooja195):

yep whats 2y+y in other words 2+1=?

OpenStudy (anonymous):

3 lol

pooja195 (pooja195):

yeah -8+5=?

OpenStudy (anonymous):

-3

pooja195 (pooja195):

3y/-3=?

OpenStudy (anonymous):

-1

pooja195 (pooja195):

\[2x+(-1)=5\] ok so now we substitute -1 into the 2nd equation to get the X value you already know that y=-1

OpenStudy (anonymous):

(3,-1) is the solution

pooja195 (pooja195):

yep good job! also... \(\huge\color{red}{WELCOME~TO~OPENSTUDY!!!}\)

OpenStudy (anonymous):

ty :D